# Learning Quadruped Locomotion using Differentiable Simulation

### v1

@@ I Introduction
$\nabla_{\theta}J(\theta)\approx\mathbb{E}[\nabla_{\theta}\log\pi_{\theta}(u_{t}|x_{t})R(\bm{\tau})],$
where it stands as a zero-order approximation of the true gradient based on sampled trajectories $\bm{\tau}$. Despite impressive real-world performance, this approach to gradient estimation is known for its high variance. Consequently, several additional strategies are required for stable training, such as using a clipped surrogate objective [31], increasing the number of samples, designing task-specific learning curriculums, employing distributed initialization techniques, and engineering suitable reward functions.
@@
In robotics, leveraging well-established knowledge about robot dynamics can enable the construction of first-order gradient estimations, which typically exhibit significantly lower variance than their zero-order counterparts and hold great potential for more stable training and faster convergence. Recently, policy training using first-order gradient has been notably advanced through differentiable simulation [8, 14, 36, 29, 37]; these works have shown promising results in reducing both the number of simulation samples and total training time compared to zero-order methods.
@@
This work set out to explore the potential of differentiable simulation in accelerating policy training within contact-rich scenarios like dynamic legged locomotion. Our goal is to establish an effective policy training framework that leverages the robot model for precise gradient computation while mitigating the difficulties posed by complex, discontinuous optimization landscapes. Moreover, we intend to evaluate the strengths and weaknesses of this approach in comparison to zero-order methods, particularly model-free policy gradients. We aim to provide useful insights for developing more efficient and robust methods that benefit from both worlds.
@@ II-C Differentiable Simulation
Differentiable simulation has recently gained momentum, thanks to the development of differentiable physics simulators [8, 5, 12] and flexible automatic differentiation tools like Jax, DiffTaichi, Pytorch [3, 13, 27]. In principle, policy training through differentiable simulation allows better convergence by replacing the zero-order gradient estimation of a stochastic objective with a more accurate estimate based on first-order gradients [33]. However, challenges such as the noisy optimization landscape and issues of exploding or vanishing gradients in long-horizon tasks render first-order gradient methods less effective.
@@ III-D Backpropagation Through Time
$\frac{\partial\mathcal{L}_{\theta}}{\partial\theta}$ $=\frac{1}{N}\sum_{k=0}^{N-1}\left(\sum_{i=1}^{k}\frac{\partial l_{i}}{\partial x_{i}}\prod_{j=i}^{k}\left(\frac{\partial x_{j}}{\partial x_{j-1}}\right)\frac{\partial x_{i}}{\partial\theta}+\frac{\partial l_{k}}{\partial u_{k}}\frac{\partial u_{k}}{\partial\theta}\right),$ (4)
@@ VI Limitations
A fundamental advantage of reinforcement learning (zero-order gradient optimization) lies in its optimization objective: RL can directly optimize non-differentiable rewards, including task-level rewards [32]. Conversely, differential simulation relies on information to be backpropagated from its differentiable loss function, making it challenging to explore novel solutions guided by task-level objectives, e.g., how to make survive differentiable?
@@ VIII Acknowledgments
This work was supported by the European Union’s Horizon Europe Research and Innovation Programme under grant agreement No. 101120732 (AUTOASSESS) and the European Research Council (ERC) under grant agreement No. 864042 (AGILEFLIGHT). The authors especially thank Yuang Zhang for his help with the software development and Seungwoo Hong for the useful discussion about rigid-body dynamics and nonlinear model predictive control. Also, the authors thank Se Hwan Jeon, Ho Jae Lee, Matthew Chignoli, Steve Heim, Elijah Stanger-Jones, and Charles Khazoom for their helpful support.

### v2

@@ I Introduction
$\nabla_{\theta}J(\theta)\approx\mathbb{E}[\nabla_{\theta}\log\pi_{\theta}(u_{t}|x_{t})R_{t}(\bm{\tau})],$
where it stands as a zeroth-order approximation of the true gradient based on sampled trajectories $\bm{\tau}$. Here, $R_{t}(\bm{\tau})$ represents the sum of future rewards. Despite impressive real-world performance, this approach to gradient estimation is known for its high variance. Consequently, several additional strategies are required for stable training, such as using a clipped surrogate objective [31], increasing the number of samples, designing task-specific learning curriculums, employing distributed initialization techniques, and engineering suitable reward functions.
@@
In robotics, leveraging well-established knowledge about robot dynamics can enable the construction of first-order gradient estimations, which typically exhibit significantly lower variance than their zeroth-order counterparts and hold great potential for more stable training and faster convergence. Recently, policy training using first-order gradient has been notably advanced through differentiable simulation [8, 14, 36, 29, 37]; these works have shown promising results in reducing both the number of simulation samples and total training time compared to zeroth-order methods.
@@
This work set out to explore the potential of differentiable simulation in accelerating policy training within contact-rich scenarios like dynamic legged locomotion. Our goal is to establish an effective policy training framework that leverages the robot model for precise gradient computation while mitigating the difficulties posed by complex, discontinuous optimization landscapes. Moreover, we intend to evaluate the strengths and weaknesses of this approach in comparison to zeroth-order methods, particularly model-free policy gradients. We aim to provide useful insights for developing more efficient and robust methods that benefit from both worlds.
@@ II-C Differentiable Simulation
Differentiable simulation has recently gained momentum, thanks to the development of differentiable physics simulators [8, 5, 12] and flexible automatic differentiation tools like Jax, DiffTaichi, Pytorch [3, 13, 27]. In principle, policy training through differentiable simulation allows better convergence by replacing the zeroth-order gradient estimation of a stochastic objective with a more accurate estimate based on first-order gradients [33]. However, challenges such as the noisy optimization landscape and issues of exploding or vanishing gradients in long-horizon tasks render first-order gradient methods less effective.
@@ III-D Backpropagation Through Time
$\frac{\partial\mathcal{L}_{\theta}}{\partial\theta}$ $=\frac{1}{N}\sum_{k=0}^{N-1}\left(\sum_{i=1}^{k}\frac{\partial l_{k}}{\partial x_{k}}\prod_{j=i}^{k}\left(\frac{\partial x_{j}}{\partial x_{j-1}}\right)\frac{\partial x_{i}}{\partial\theta}+\frac{\partial l_{k}}{\partial u_{k}}\frac{\partial u_{k}}{\partial\theta}\right),$ (4)
@@ VI Limitations
A fundamental advantage of reinforcement learning (zeroth-order gradient optimization) lies in its optimization objective: RL can directly optimize non-differentiable rewards, including task-level rewards [32]. Conversely, differential simulation relies on information to be backpropagated from its differentiable loss function, making it challenging to explore novel solutions guided by task-level objectives, e.g., how to make survive differentiable?
@@ VIII Acknowledgments
This work was supported by the European Union’s Horizon Europe Research and Innovation Programme under grant agreement No. 101120732 (AUTOASSESS) and the European Research Council (ERC) under grant agreement No. 864042 (AGILEFLIGHT). The authors especially thank Yuang Zhang for his help with the software development, Seungwoo Hong for the useful discussion about rigid-body dynamics and nonlinear model predictive control, and Nico Messikommer for the valuable feedback. Also, the authors thank Se Hwan Jeon, Ho Jae Lee, Matthew Chignoli, Steve Heim, Elijah Stanger-Jones, and Charles Khazoom for their helpful support.